OpenStudy (anonymous):

A number cube with the numbers 1, 2, 3, 4, 5, and 6 is rolled 1000 times. About how many times would it be expected that a number greater than 4 is rolled? A. 667 B. 167 C. 500 D. 333

OpenStudy (smored):

there are two (5 and 6) numbers that are greater than 4. that is 2 out of six, so the percentage is \[\frac{2}{6}=\frac{1}{3}\]which is about 0.333 (33.33%). 33.33% of 1000 is \[1000(0.333)=333\]

OpenStudy (nirmalnema):

2/6 = x/1000 x=333